OpenStudy (anonymous):

Use the given information to find the minimum sample size required to estimate an unknown population mean μ. Margin of error: $126, confidence level: 99%, σ = $512

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)

OpenStudy (yanasidlinskiy):

Margin of Error (half of confidence interval) 136 The margin of error is defined as the "radius" (or half the width) of a confidence interval for a particular statistic. Level of Confidence \(\huge{\rightarrow}\) 99 s = Sample standard deviation \(\huge{\rightarrow}\) 512

OpenStudy (yanasidlinskiy):

'z critical value' from Look-up Table for 99% What's the #?

OpenStudy (anonymous):

2.6

OpenStudy (yanasidlinskiy):

Ok. When you have 2.6, significant digits 2 Margin of Error = ('z critical value') * \[\frac{ s }{ \sqrt{n} }\] n = .....?

OpenStudy (anonymous):

How do you get n?

OpenStudy (yanasidlinskiy):

2.6 * \[\frac{ s}{ \sqrt{n} }\]

OpenStudy (yanasidlinskiy):

Mulitply.